OpenStudy (anonymous):

Solve the sytem by elimination: 3x-y=7 2x+y=3

OpenStudy (anonymous):

if you add the two equations together, you will eliminate the y variable: 3x - y + 2x + y = 7 + 3, which gives you 5x = 10 or x = 2. now you can substitute the 2 in either equation and find out what the y is

OpenStudy (anonymous):

Ah, ok. Thanks for the help.

OpenStudy (anonymous):

no prob